Find the solution of the recurrence relation

an = 4an−1 − 3an−2 + 2


n + n + 3 , a0 = 1 and a1 = 4


1
Expert's answer
2021-12-15T17:01:27-0500

Let us find the solution of the recurrence relation "a_n = 4a_{n\u22121} \u2212 3a_{n\u22122} + 2^n + n + 3" with

"a_0 = 1" and "a_1 = 4."

The characteristic equation "k^2-4k+3=0" of the homogeous recurrence relation "a_n -4a_{n\u22121} + 3a_{n\u22122}=0" is equivalent to "(k-1)(k-3)=0," and hence has the solutions "k_1=1"

and "k_2=3." Therefore, the general solution of the relation "a_n = 4a_{n\u22121} \u2212 3a_{n\u22122} + 2^n + n + 3" is of the form "a_n=c_1+c_23^n+b_n^p," where "b_n^p=a2^n+n(bn+c)=a2^n+bn^2+cn."

It follows that

"a2^n+bn^2+cn"

"=4(a2^{n-1}+b(n-1)^2+c(n-1))-3(a2^{n-2}+b(n-2)^2+c(n-2))+2^n+n+3"

"=4a2^{n-1}+4b(n^2-2n+1)+4c(n-1)-3a2^{n-2}-3b(n^2-4n+4)-3c(n-2)+2^n+n+3"

"=4a2^{n-1}-3a2^{n-2}+2^n+bn^2+(4b+c+1)n+(-8b+2c+3)"

"=(5a+4)2^{n-2}+bn^2+(4b+c+1)n+(-8b+2c+3)."

It follows that "4a=5a+4,\ c=4b+c+1" and "-8b+2c+3=0."

Therefore, "a=-4,\ b=-\frac{1}4,\ c=\frac{1}2(8b-3)=\frac{1}2(-2-3)=-\frac{5}2."

We conclude that the general solution of the relation "a_n = 4a_{n\u22121} \u2212 3a_{n\u22122} + 2^n + n + 3" is of the form "a_n=c_1+c_23^n-4\cdot2^n-\frac{1}4n^2-\frac{5}2n."

Since "a_0 = 1" and "a_1 = 4," we get

"1=a_0=c_1+c_2-4" and "4=a_1=c_1+3c_2-8-\frac{1}4-\frac{5}2=c_1+3c_2-\frac{43}4."

Therefore, "c_1+c_2=5" and "c_1+3c_2=\frac{59}4." It follows that "c_2=\frac{39}8" and "c_1=\frac{1}8."


Consequently, the general solution of the relation "a_n = 4a_{n\u22121} \u2212 3a_{n\u22122} + 2^n + n + 3" is the following:

"a_n=\frac{39}8+\frac{1}83^n-2^{n+2}-\frac{1}4n^2-\frac{5}2n."